OpenStudy (anonymous):

Convert 12.0 kilograms into grams.

OpenStudy (anonymous):

1 kg = 1000g

OpenStudy (anonymous):

1 kg = 1000g 12 kg = 12*1000 g = ?

hero (hero):

12kg = xg Divide both sides by g to get: 12k = x Change k to 1000 to get 12*(1000) = x